Solve the recurrence relation a_n=-3a_(n-1)-3a_(n-2)-a_(n-3) with a_0=5,a_1=-9,and a_2=15.
1
Expert's answer
2020-06-02T19:49:56-0400

Solution.

"a_n=-3a_{n-1}-3a_{n-2}-a_{n-3}, a_0=5, a_1=-9, a_2=15;"

Let "a_n=r^3, a_{n-1}=r^2, a_{n-2}=r, a_{n-3}=1,"then

"r^3=-3r^2-3r-1;"

"r^3+3r^2+3r+1=0";

"(r+1)^3=0;"

"r+1=0;"

"r=-1;"

"a_n=\alpha_1r^n+\alpha_2nr^n+ … +\alpha_kn^k-1r^n;"

So r=-1, then "a_n=\alpha_1(-1)^n+\alpha_2n(-1)^n+\alpha_3n^2(-1)^n;"

Substitute the value and find "\alpha_1,\alpha_2,\alpha_3:"

"5=\alpha_1(-1)^0, \alpha_1=5;"

"-9=\alpha_1(-1)^1+\alpha_2(-1)^1+\alpha_3(-1)^1=-\alpha_1-\alpha_2-\alpha_3;"

"15=\alpha_1(-1)^2+\alpha_2\sdot2\sdot(-1)^2+\alpha_3\sdot4\sdot(-1)^2=\alpha_1+2\alpha_2+4\alpha_3;"

"-9=-5-\alpha_2-\alpha_3;" | "\sdot2"

"15=5+2\alpha_2+4\alpha_3;"

"-18=-10-2\alpha_2-2\alpha_3;"

"15=5+2\alpha_2+4\alpha_3;"

____________________________

"-3=-5+2\alpha_3;"

"2=2\alpha_3;"

"\alpha_3=1;"

"15=5+2\alpha_2+4;"

"6=2\alpha_2;"

"\alpha_2=3;"

"a_n=5(-1)^n+3n(-1)^n+1n^2(-1)^n=(-1)^n(5+3n+n^2);"

Answer: "a_n=(-1)^n(5+3n+n^2)."